OpenStudy (anonymous):

A coin was tossed 4 times, find the chance of getting 2 heads

OpenStudy (anonymous):

the chance for 1 head 1/4 for 4 tosses. for 2 heads its gonna be 1/8

OpenStudy (kropot72):

The binomial distribution gives the probability of 2 heads in 4 tosses as \[P(2\ heads)=\left(\begin{matrix}4 \\ 2\end{matrix}\right)0.5^{2}0.5^{2}=\frac{6}{16}=\frac{3}{8}\]